my question is find the equation of a line passing through the point (5,-3) that is perpendicular to the line 3x + 4y=24. use slope intercept formula.
thanks a lot

Answer by dolly(163)

Hi,
The given line is 3x + 4y = 24
==> 4y = - 3x + 24
==> y = - (3/4)x + 6 [ dividing by 4]
This is the slope intercept form where the slope is -(3/4)
So the slope of the required line = (4/3) [ as the product of the slopes of perpendicular lines is -1]
This line passes through (5, -3)
So using the slope point form the equation is:
y - (-3) = (4/3)(x-5)
==> y + 3 = (4/3)(x-5)
==> 3(y+3) = 4(x-5)
==> 3y + 9 = 4x - 20
==> 3y = 4x - 20 - 9
==> 3y = 4x - 29
Thus the equation of the required line is 3y = 4x - 29 or 4x - 3y = 29
good luck!!!
dolly